### v3

Abstract
With the growing application of computer vision in agriculture, image analysis has become essential for tasks such as crop health monitoring and pest detection. However, significant domain shifts caused by environmental variations, different crop types, and diverse data acquisition methods hinder model generalization across regions, seasons, and complex agricultural settings. This paper investigates how Domain Adaptation (DA) techniques can address these challenges by improving cross-domain transferability in agricultural image analysis. Given the limited availability of labeled data, weak model adaptability, and dynamic field conditions, DA has emerged as a promising solution. The review systematically summarizes recent advances in DA for agricultural imagery, focusing on applications such as crop health monitoring, pest detection, and fruit recognition, where DA methods have enhanced performance across diverse domains. DA approaches are categorized into shallow and deep learning methods, including supervised, semi-supervised, and unsupervised strategies, with particular attention to adversarial learning-based techniques that have demonstrated strong potential in complex scenarios. In addition, the paper reviews key public agricultural image datasets, evaluating their strengths and limitations in DA research. Overall, this work offers a comprehensive framework and critical insights to guide future research and development of domain adaptation in agricultural vision tasks.
keywords:
Domain Adaptation, Deep Learning, Agricultural Image Analysis, Cross-Domain Transfer, Adversarial Learning
1 Introduction
In agricultural image analysis, the high heterogeneity of data acquisition environments and the extreme scarcity of annotated data severely limit the generalization capability of models across different application scenarios. Among these challenges, the domain shift problem—referring to distribution discrepancies between source and target data—has emerged as a core obstacle hindering the deployment of agricultural intelligent vision systems [1, 2]. To address this, domain adaptation (DA) techniques have been widely introduced into agricultural image analysis in recent years, aiming to enhance model robustness and transferability across cross-regional, cross-temporal, and cross-device data.
Agricultural image data acquisition typically relies on multiple sensors (e.g., RGB cameras, near-infrared cameras, hyperspectral sensors) and is susceptible to complex environmental factors such as illumination variations, seasonal changes, crop varieties, and soil backgrounds. These differences lead to significant distribution shifts for the same agricultural task across datasets, causing models trained under conventional paradigms to suffer drastic performance degradation in new environments [3]. For instance, models trained in laboratories or demonstration farms often exhibit sharp accuracy drops when deployed in actual fields, different regions, or across growing seasons.
Moreover, annotating agricultural images requires expertise from agronomists and plant protection specialists, resulting in high costs and limited scalability. Compared to general computer vision tasks, the availability of high-quality annotated datasets in agriculture remains extremely scarce [4]. Consequently, directly transferring pre-trained models from natural image domains proves ineffective, underscoring the necessity of domain adaptation to reduce annotation dependency and enhance cross-domain adaptability. The core of domain adaptation lies in minimizing feature distribution discrepancies between source and target domains, enabling models to achieve comparable performance on the target domain.
Recently, DA techniques have demonstrated significant efficacy in key tasks such as crop disease/pest detection, hyperspectral image processing, crop growth monitoring, and precision farming management. Particularly under unsupervised or weakly supervised conditions, their ability to model data distribution shifts positions them as a critical pathway toward scalable agricultural visual intelligence. Although numerous surveys on transfer learning and domain adaptation exist in natural image analysis [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16], systematic reviews tailored to agricultural image analysis remain insufficient. Building upon this, this survey aims to systematically review and summarize the current state of domain adaptation methods in agricultural image analysis. We categorize existing methods into two major classes (shallow and deep learning models) and discuss their supervised, semi-supervised, and unsupervised strategies. Particular focus is placed on adversarial learning-based adaptation methods and their responses to challenges such as seasonal variation, multi-source fusion, and annotation scarcity, ultimately offering a methodological reference for future intelligent agricultural systems.
2 Background
2.1 Domain Shift in Agricultural Image Analysis
In agricultural image analysis, domain shift refers to the discrepancy in data distribution between the training dataset (source domain) and the test dataset (target domain). This shift can lead models to perform well on the source domain but fail to generalize effectively to the target domain, thereby significantly degrading overall model performance. Agricultural images often exhibit substantial distributional differences due to a variety of factors, including environmental variations, crop types, and imaging conditions. Specifically, images captured in different geographic regions or under varying environmental settings—such as changes in lighting, climate, seasonal dynamics, crop varieties, and camera viewpoints—exacerbate domain shift and present serious challenges to model transferability.
As shown in Figure 1, numerous studies have explored the use of domain adaptation (DA) techniques to address these challenges in agricultural vision tasks. DA methods aim to mitigate domain shift by aligning feature distributions between source and target domains, enabling more robust and generalizable model performance.
Figure 1: Illustration of source and target data: (left) the original feature distribution shows misalignment due to domain shift; (right) improved feature alignment after applying domain adaptation methods.
For example, in crop classification tasks, images collected from different regions often exhibit marked heterogeneity in appearance. Figure 2 illustrates this phenomenon using aerial scenes from Ghaziabad, India, captured during the rainy and dry seasons. The images clearly reflect substantial environmental differences, such as lighting and climate, that introduce distributional discrepancies between source and target domains. These shifts significantly impact the accuracy and robustness of key agricultural vision tasks, including crop monitoring and pest detection, particularly in cross-regional, cross-seasonal, and cross-device scenarios.
As the demand for scalable agricultural image analysis continues to grow, domain adaptation has emerged as a critical technology for improving cross-domain generalization. By reducing distributional gaps between domains, DA techniques enhance the reliability of agricultural models across diverse settings, ultimately supporting more accurate, consistent, and transferable performance in real-world agricultural applications.
Figure 2: Aerial scenes from Ghaziabad, India, captured during rainy and dry seasons, illustrating domain shift due to seasonal variation.
2.2 Domain Adaptation and Transfer Learning
In a typical transfer learning (TL) framework, two fundamental concepts are involved: domain and task. A domain refers to a specific feature space along with its marginal probability distribution, while a task is defined by a label space and an associated predictive function. The goal of transfer learning is to leverage knowledge learned from a source task $T_{A}$ in domain $A$ and apply it to a target task $T_{B}$ in domain $B$. Notably, transfer learning allows for either the domain or the task—or both—to differ between the source and target settings.
Domain Adaptation (DA) is a specialized and widely studied subfield of transfer learning. In DA, the feature space and task remain consistent across domains, but the marginal distributions differ, i.e., $P_{S}(X)\neq P_{T}(X)$. Formally, let $X\times Y$ denote the joint feature and label space. The source domain $S$ and target domain $T$ are defined over this space, but their respective distributions $P_{S}$ and $P_{T}$ differ. Assume there are $n_{S}$ labeled samples in the source domain, $D_{S}=\{(x_{i}^{S},y_{i}^{S})\}_{i=1}^{n_{S}}$, and $n_{T}$ samples (either labeled or unlabeled) in the target domain, $D_{T}=\{x_{j}^{T}\}_{j=1}^{n_{T}}$.
The objective of domain adaptation is to transfer knowledge from the source domain $S$ to improve model performance on the target domain $T$ for a shared task. This is typically achieved by aligning the distributions of the two domains—either explicitly or implici
2.3 Domain Adaptation Problem Setup
In the agricultural domain, domain adaptation (DA) methods face unique challenges stemming from pronounced discrepancies between source and target domains. These discrepancies are particularly evident in three dimensions: spatial distribution, temporal sequences, and modality diversity. Agricultural image data are typically collected from heterogeneous platforms, including satellite remote sensing, unmanned aerial vehicles (UAVs), and ground-based sensors. Due to variations in climate conditions, crop types, and sensor characteristics across different environments, substantial distributional heterogeneity arises among agricultural datasets. To address this, domain adaptation techniques have become essential tools for enhancing model generalization by bridging distribution gaps across domains. Depending on the availability of labeled data in the target domain, DA methods are commonly categorized into supervised, semi-supervised, and unsupervised approaches, with unsupervised domain adaptation (UDA) drawing particular attention due to the scarcity and high cost of annotations in agricultural settings.
Early efforts in domain adaptation primarily focused on aligning the data distributions between source and target domains through three main technical pathways: instance-based alignment (e.g., sample reweighting), feature space transformation (e.g., subspace learning), and classifier parameter adaptation. These traditional methods typically rely on hand-crafted features and statistical alignment techniques to reduce domain discrepancy. For instance, instance-based approaches adjust the contribution of each source sample to better match the target distribution; feature-based methods seek a shared subspace through linear projections; and classifier-based methods improve cross-domain performance by regularizing classifier parameters. These techniques, collectively referred to as shallow domain adaptation (Shallow DA), offer advantages such as low computational complexity, high training efficiency, and strong interpretability. Consequently, they are particularly well-suited for deployment in resource-constrained agricultural scenarios.
The emergence of deep learning significantly reshaped the domain adaptation landscape. In 2014, Girshick et al. [17] introduced the Region-based Convolutional Neural Network (R-CNN), pioneering the use of convolutional neural networks (CNNs) for object detection. Building on this, Fast R-CNN [18] and later Faster R-CNN [19] integrated region proposal, feature extraction, and classification into a unified, end-to-end architecture, vastly improving both accuracy and efficiency. In the agricultural context, Fuentes et al. [20] applied the Faster R-CNN framework with VGGNet/ResNet backbones to detect and localize tomato diseases and pests, achieving a mean Average Precision (mAP) of 85.98% across 10 classes. This success exemplifies the widespread adoption of deep learning in agricultural image analysis. Owing to their reliance on deep feature extraction and complex architectures, these approaches fall under the umbrella of deep domain adaptation (Deep DA).
In summary, domain adaptation methods in agricultural image analysis can be systematically divided into two major paradigms: shallow and deep adaptation. Shallow DA includes classical techniques based on instance reweighting, feature alignment, and classifier adjustment, while Deep DA encompasses modern methods that exploit deep neural networks and can be further categorized into unsupervised, semi-supervised, and supervised variants depending on target label availability (see Table I). These two paradigms offer complementary strengths in terms of computational cost, scalability, and domain transferability, together forming a comprehensive framework for addressing cross-domain challenges in agricultural visual perception.
Table 1: Taxonomy of Domain Adaptation Methods in Agricultural Image Analysis
CATEGORY SUBCATEGORIES TYPICAL EXAMPLES
Shallow DA Instance-based DIAFAN-TL, RHM, HSI, CSSPL
Feature-based SA, TSSA, ECMDCM, CropSTGAN, MultiCropGAN, Spatial-Invariant Features, Invariant Features in RKHS, Sensor-driven Hierarchical DA, PSO-TrAdaBoost, ESMSL, HSL-GM
Classifier-based MLCA, BCC, SSGF, MDAF, WRF, CDELM, BHC, DASVM, SD-AL, AMKFL, VSV, EasyTL
Deep DA Semi-supervised learning MAML, WheatSeedBelt, SSDA-WheatHead, SSDA-WheatSeg, TDA-YOLO, SCDAL, CDADA
Unsupervised learning MSFF, GAN-DA, Standard GAN, DeepDA-Net, U-DA-Net, TriADA, MaxEnt-DA, DAE-DANN, Self-Attn-DA, ADANN, OpenDA, AdaptSegNet, ADVENT, BDL, DFENet, TSAN, DAN, WDGRL, CORAL, Transformer-UDA, MRAN, WPS-DSA, NCADA, TCANet, DDA-Net, TDDA, TST net, AMF-FSL, MSUN, MSCN, AMRAN, DJDANs, CLA
3 Shallow DA Methods
This section reviews shallow domain adaptation (DA) methods that rely on manually engineered features and traditional machine learning algorithms in the context of agricultural data analysis. These methods are typically categorized into three main strategies: (1) instance weighting, (2) feature transformation, and (3) classifier-based adaptation.
3.1 Instance Weighting
Instance weighting is a widely adopted strategy in shallow DA, particularly in agricultural image analysis. These methods aim to adjust the marginal distributions of the source and target samples to reduce domain discrepancies. Let $p_{t}(x)$ and $p_{s}(x)$ denote the marginal density functions of the target and source samples, respectively. The importance weight $w(x)$ is defined as:
$w(x)=\frac{p_{t}(x)}{p_{s}(x)}$ (1)
In this approach, source domain samples are weighted according to their similarity or relevance to target domain samples. For example, crop observation data from geographically adjacent regions might be assigned higher weights to better reflect the target domain distribution. Training a model—such as a classifier or regressor—on these reweighted samples can significantly mitigate the distributional gap between domains. As illustrated in Figure 3, instance weighting visibly reduces the divergence between source and target domains.
Figure 3: Instance weighting to alleviate domain shift. (a) Source domain after feature matching. (b) Target domain after feature matching. (c) Source domain with joint feature matching and instance weighting, where unfilled markers indicate less relevant, low-weight instances. Image adapted from Longet et al. [21].
Molina-Cabanillas et al. [22] introduced DIAFAN-TL (Domain-Invariant Agricultural Features Adaptation Network for Transfer Learning), an instance-weighted learning algorithm for phenological prediction in olive tree cultivars. This method dynamically assigns weights to training samples based on inter-domain relevance, thus enhancing information diversity while reducing sample inconsistency.
Yaras et al. [23] proposed Random Histogram Matching (RHM), which leverages data augmentation to improve satellite image domain adaptation. By modeling sensor differences and environmental variations as nonlinear pixel transformations, and applying deep neural networks to augment data, this method boosts robustness to domain shifts.
For further framework optimization, Cui et al. [24] developed the Iterative Weighted Active Transfer Learning (IWATL) approach for hyperspectral image (HSI) classification. Through a dual-evaluation mechanism, it adaptively updates source sample weights, significantly enhancing classification performance. Li et al. [25] proposed the Cost-Sensitive Self-Paced Learning (CSSPL) framework, which integrates hybrid weight regularization for selecting multi-temporal samples, addressing seasonal and temporal variances in agricultural monitoring.
In summary, instance-based shallow DA methods typically follow a two-stage optimization process: (1) adjusting the sample distribution via reweighting or selection, and (2) integrating the adapted data into a robust classifier. These compound strategies are particularly effective in agricultural remote sensing, where data heterogeneity is prevalent, and have demonstrated clear improvements in transferability and generalization.
3.2 Feature Transformation
Feature-based domain adaptation methods aim to map data from the source and target domains into a shared, domain-invariant feature representation space. This is typically achieved through techniques such as subspace alignment, manifold learning, and low-rank representation. The primary objective of these methods is to reduce the distribution discrepancy between domains during the process of joint feature extraction [26].
As illustrated in Figure 4, these methods first construct a common feature space that captures domain-invariant properties. Subsequently, a classifier is trained using the transformed source domain features and corresponding labels, enabling the prediction of target domain samples with improved accuracy. Implementation strategies generally fall into two categories: subspace-based projection reconstruction and nonlinear transformation-based feature re-encoding.
The overarching goal is to align the statistical properties of features across domains by transforming the feature space, thereby enhancing the generalization ability of the model on the target domain. This approach is particularly effective in agricultural scenarios characterized by domain shifts induced by seasonal, geographical, or sensor-based variations.
Figure 4: Feature transformation for domain adaptation. Each color represents a distinct domain. (a) Feature distributions of two source domains and one target domain before transformation. (b) Aligned feature distributions after transformation. Image adapted from Wang et al. [27].
3.2.1 Subspace-Based Adaptation
Subspace-based domain adaptation (DA) methods typically project source and target samples into separate subspaces using subspace learning or dimensionality reduction techniques, and then align the subspaces. The Subspace Alignment (SA) method proposed by Fernando et al. [28] has been extended to multispectral agricultural image classification tasks. By employing Principal Component Analysis (PCA), this method aligns crop spectral feature subspaces across different seasons, thereby alleviating domain shifts caused by lighting and phenological variations. The SA process is illustrated in Figure 5.
Figure 5: Illustration of Subspace Alignment.
As shown, the source domain samples $X_{S}\in\mathbb{R}^{n_{s}\times D}$ contain $n_{s}$ samples with $D$-dimensional features, and the target domain samples $X_{T}\in\mathbb{R}^{n_{t}\times D}$ share the same dimensionality. PCA is applied to obtain the principal components of source and target domains, and a linear transformation matrix $M$ is learned for subspace alignment.
The core formula for subspace alignment is:
[…]
where $P_{s}$ and $P_{t}$ are the projection matrices (e.g., PCA bases) of the source and target subspaces, respectively, and $\|\cdot\|_{F}^{2}$ denotes the Frobenius norm used to measure domain discrepancy.
Sun et al. [29] applied SA to cross-view scene classification using Partial Least Squares (PLS) to construct a discriminative subspace shared by both domains, effectively capturing common spectral features across different sensors. They further proposed the Transfer Sparse Subspace Analysis (TSSA) method [30], which incorporates sparse subspace clustering and Maximum Mean Discrepancy (MMD) minimization to retain self-expressive properties of data across maize growth stages.
Weilandt et al. [31] introduced the Early Crop Mapping based on Dynamic Clustering Method (ECMDCM), leveraging time-series NDVI and EVI data for improved accuracy in early-stage crop classification. Li et al. [32] proposed CropSTGAN, which includes a domain mapper to adapt spatial-temporal discrepancies in crop mapping, while Wang et al. [33] developed MultiCropGAN, incorporating identity loss to preserve essential features during label space adaptation, with validation in North American farmlands.
Adversarial learning techniques based on category alignment offer another avenue. Takahashi et al. [34] proposed a category-guided feature alignment framework, which maps features from both domains into a shared semantic space using adversarial training. Without requiring target domain labels, it leverages source domain categories to enhance domain-invariant learning, improving generalization across crop types and acquisition conditions.
Feature-invariant methods can be viewed as a specific class of subspace adaptation. These methods aim to identify features that remain stable across domains. Bruzzone et al. [35] introduced a multi-objective optimization approach for selecting spatially invariant features to classify non-overlapping scenes, balancing discriminability and spatial robustness. Invariant features may also be extracted in Reproducing Kernel Hilbert Spaces (RKHS) [36]. Paris et al. [37] proposed a sensor-driven hierarchical DA strategy using such features, and Yan et al. [38] introduced a PSO-based TrAdaBoost variant for optimal subspace selection in cross-domain classification.
To address challenges in hyperspectral image classification, such as label-feature inconsistency and computational inefficiency in similarity learning, Yi [39] proposed a subspace learning framework based on set similarity metrics. Compared with traditional methods like Linear Discriminant Analysis (LDA), this approach improves classification accuracy by 8%–15% and reduces training time by 40%–60% on datasets such as Indian Pines and Salinas Valley.
Finally, Banerjee [40] proposed a hierarchical subspace alignment method using a semantic-driven binary tree and Grassmannian manifold alignment to address multi-temporal distribution mismatches due to atmospheric, sensor, and labeling differences. It outperforms traditional techniques (e.g., Global GFK, SA) by 5%–6% in accuracy and substantially reduces misclassification in semantically similar crop categories, offering a scalable solution for dynamic agricultural monitoring and environmental analysis.
3.2.2 Transformation-Based Adaptation
Transformation-based Domain Adaptation (DA) methods aim to reduce domain discrepancies by directly modifying the distribution of data or features, rather than projecting them into aligned subspaces. These approaches typically apply mapping functions to transform data into a shared representation space or perform direct data-level transformations, thereby aligning the statistical properties of source and target domains.
Aptoula et al. [41] proposed a deep learning–based framework for domain-adaptive classification of crops and weeds, addressing distributional differences caused by variations in lighting, soil properties, and plant phenology. This method embodies the essence of transformation-based DA by projecting agricultural remote sensing data into a new feature space that minimizes both statistical and geometric distributional differences across domains, while preserving structural consistency. These methods have demonstrated substantial utility in agricultural remote sensing and crop monitoring applications (see Figure 6).
Figure 6: Illustration of Transformation-Based Adaptation.
The transformation process typically relies on metrics such as Maximum Mean Discrepancy (MMD), Kullback-Leibler (KL) divergence, or Bregman divergence to quantify inter-domain differences. Among these, MMD is widely favored for agricultural data due to its robustness in high-dimensional settings, and is defined as:
$\text{MMD}(X_{s},X_{t})=\left\|\frac{1}{n_{s}}\sum_{\mathbf{x}_{i}\in X_{s}}\phi(\mathbf{x}_{i})-\frac{1}{n_{t}}\sum_{\mathbf{x}_{j}\in X_{t}}\phi(\mathbf{x}_{j})\right\|_{F}^{2}$ (3)
where $\phi(\cdot)$ denotes the mapping to a Reproducing Kernel Hilbert Space (RKHS), and $X_{s}$, $X_{t}$ are the source and target domain datasets, respectively.
The effectiveness of different kernel functions in modeling high-dimensional agricultural data (e.g., hyperspectral imagery) lies in their ability to capture nonlinear and complex feature relationships:
  • Linear Kernel: Suitable for linearly separable data but often inadequate for hyperspectral imagery due to the data’s inherent nonlinearity.
  • Radial Basis Function (RBF) Kernel: Ideal for hyperspectral data as it captures local nonlinear variations across spectral bands, offering a balance between expressiveness and computational efficiency.
  • Polynomial Kernel: Can model nonlinear feature relationships, especially when inter-band correlations exist; however, high-degree polynomials may incur overfitting and computational costs.
  • Sigmoid Kernel: Mimics neural activations but is rarely used in high-dimensional agricultural data due to training instability and poor scalability.
In general, the RBF kernel remains the preferred choice for hyperspectral and other complex agricultural data types due to its robust performance in capturing nonlinear structures.
Pan et al. [42] introduced Transfer Component Analysis (TCA), which projects source and target data into a shared RKHS and minimizes domain discrepancy via MMD. The TCA objective function is:
$\min\operatorname{tr}(W^{\top}KLK^{\top}W)+\mu\operatorname{tr}(W^{\top}W)$ (4)
where $K$ is the kernel matrix, $L$ is the MMD matrix, $W$ is the transformation matrix, and $\mu$ is the regularization parameter. TCA’s adaptability to varied growth stages and its low annotation dependency make it well-suited for tasks such as pest detection, yield estimation, and resource allocation.
Canonical Correlation Analysis (CCA) and its derivatives also provide effective solutions for heterogeneous DA in agricultural contexts. These transformation-based methods, by innovatively modeling domain shifts and tailoring solutions to agricultural data characteristics, contribute significantly to improving model robustness and cross-domain generalization in smart farming systems.
3.3 Classifier-Based Adaptation
Classifier-based domain adaptation (DA) methods refine classifier parameters trained on the source domain by incorporating unlabeled or sparsely labeled samples from the target domain. These approaches have proven particularly valuable in agricultural scenarios characterized by diverse environments, seasonal variability, and limited annotations.
In agriculture, domain shifts often arise due to seasonal changes, varying crop phenotypes, and different imaging conditions. To address these, Bruzzone et al.[43, 44, 45, 46] introduced a maximum likelihood classifier adaptation method, which iteratively updates classifier parameters to better reflect the statistical properties of new-season imagery. Similarly, an improved Bayesian cascade classifier[44] enhanced the robustness of cross-regional crop-type recognition. Zhong et al.[47] integrated spectral-spatial guided filtering with adaptive classifier updating, significantly boosting the detection accuracy of crop diseases in heterogeneous field conditions.
Given the complexity of cross-domain agricultural data, ensemble learning has gained traction as an effective strategy to increase robustness. By aggregating predictions from multiple base classifiers, ensemble techniques mitigate overfitting and model bias. Wei et al.[48] proposed a Multi-Domain Adaptation Fusion (MDAF) method that improves farmland classification accuracy across different countries through classifier ensembles, addressing spectral variability.
Weighted classifier transfer approaches—especially those based on Random Forests—are also extensively employed in agriculture. Zhang et al.[49] utilized weighted Random Forests to merge agricultural data from distinct geographic regions, enhancing the performance of cross-regional crop disease detection. Xu et al.[50] extended Extreme Learning Machine (ELM) transfer with kernel alignment to predict crop yields under varying climate conditions, further exemplifying the utility of classifier-based adaptation in agriculture.
To cope with the high cost of labeling agricultural imagery, semi-supervised learning (SSL) and active learning (AL) have become increasingly prominent. Rajan et al.[51] proposed a Binary Hierarchical Classifier (BHC) integrating SSL for continuous model updates in cross-year pest monitoring. Additionally, enhancements to the classic Domain Adaptation SVM (DASVM) have enabled improved generalization in cross-sensor crop classification with minimal labeled target samples[52].
Figure 7: Flowchart of an Active Learning (AL) method for domain adaptation.
Figure 7 illustrates a typical AL process. Initially, a classifier is trained on the source domain and used to label the target domain. Samples with the highest prediction uncertainty are selected for manual labeling and added back to the training set, refining the classifier iteratively. AL is highly effective in precision agriculture due to its ability to reduce annotation demands.
Deng et al.[53] developed an Active Multi-Kernel Domain Adaptation (AMKDA) method combining AL with multi-kernel learning, substantially reducing the need for labeled hyperspectral data. Kalita et al.[54] introduced a standard deviation (SD)-based AL strategy that identifies the most informative samples, while Saboori et al.[55] proposed Active Multi-Kernel Fredholm Learning (AMKFL), which uses Fredholm regularization to guide annotation and improve classification.
To address morphological and scale variations in agricultural imagery, innovative classifiers have been introduced. Izquierdo-Verdiguier et al.[56] proposed the Virtual Support Vector (VSV) classifier, which incorporates rotation and scale invariance, achieving up to a 12% improvement in UAV-based farmland classification. Wang et al.[57] developed EasyTL, a non-parametric transfer method suitable for edge computing, enabling real-time crop status recognition with high adaptability in resource-limited agricultural environments.
3.4 Summary of Shallow DA Methods
Shallow Domain Adaptation (DA) methods primarily rely on traditional machine learning algorithms and handcrafted feature representations to mitigate distributional discrepancies between the source and target domains. These approaches can be broadly categorized into three types: instance weighting, feature transformation, and classifier-based adaptation. Instance-based methods reweight source domain samples based on their relevance to the target domain, thereby reducing marginal distribution shifts. Feature transformation techniques aim to project both domains into a shared, domain-invariant subspace to achieve feature-level alignment. Classifier-based adaptation methods focus on modifying classifier parameters to better generalize to target domain data.
The key advantages of shallow DA methods include their low computational complexity, ease of implementation, and reduced dependency on large-scale computational resources. These traits make shallow approaches particularly suitable for agricultural applications in resource-limited environments, such as edge computing systems or developing regions.
However, shallow methods exhibit notable limitations when applied to high-dimensional, structurally complex, and large-scale agricultural datasets. Due to their limited capacity for hierarchical feature extraction, they often struggle to model nonlinear relationships and contextual dependencies inherent in real-world agricultural scenarios. In cases involving substantial domain shifts—such as cross-regional crop mapping or multi-seasonal disease detection—shallow methods are generally inadequate for capturing deep semantic correlations, which compromises their generalization performance.
These limitations have catalyzed the adoption of deep learning–based domain adaptation techniques. The ability of deep models to learn rich, multi-level feature representations and capture complex domain-specific patterns has made them a powerful alternative to traditional methods, offering more robust and scalable solutions for cross-domain agricultural image analysis.
4 Deep Domain Adaptation (DA) Methods
With the rapid advancement of deep learning technologies, domain adaptation (DA) methods based on deep neural networks have demonstrated notable advantages in agricultural image analysis. These methods automatically extract domain-invariant features through end-to-end feature learning and distribution alignment, effectively mitigating domain shift problems caused by variations in sensors, environmental conditions, and crop growth stages. Depending on the availability of labeled data in the target domain, deep DA methods are typically classified into supervised, semi-supervised, and unsupervised approaches. This section provides a systematic review of their core principles and recent developments in agricultural applications.
4.1 Supervised Deep DA
Supervised deep DA methods assume access to a limited amount of labeled data in the target domain. These methods jointly optimize the feature extractor and classifier by minimizing the combination of source domain classification loss and cross-domain alignment loss. The primary idea is to construct a deep neural network that integrates target domain label supervision into the domain alignment process. A typical supervised deep DA model consists of three modules: a feature extractor, a classifier, and a domain discrepancy measurement component. The joint loss function is formulated as:
$L=L_{cls}(X_{s},Y_{s})+\lambda\cdot L_{align}(X_{s},X_{t})$ (5)
Here, $L_{cls}(X_{s},Y_{s})$ denotes the classification loss on the source domain, $L_{align}(X_{s},X_{t})$ measures the domain discrepancy (e.g., via MMD, adversarial loss, or covariance alignment), and $\lambda$ is a trade-off parameter.
In agricultural scenarios, supervised deep DA is widely used for cross-sensor data fusion and multi-temporal crop monitoring. For instance, Takahashi et al. [34] proposed a category-aware adversarial learning framework that utilizes source domain class labels to guide feature space alignment, thereby improving generalization on target agricultural datasets collected from heterogeneous imaging devices (see the Figure 8).
Figure 8: Structure of the method by Takahashi et al. [34].
This approach comprises three main stages: image generation, transfer learning, and domain adaptation based on deep neural networks (DNNs). First, a Class Propagation Network (CPN) generates a class map from target images, simulating realistic agricultural image distributions. Second, transfer learning is performed on both real images ($I$) and synthetic images ($W$), producing two transfer models, TL-I and TL-W. Finally, deep networks implement domain adaptation using two modules: DA-W2I (synthetic-to-real adaptation) and DA-I2W (real-to-synthetic adaptation).
As shown in Figure 8, the system architecture includes five components: CPN for image generation, TL-I and TL-W for transfer learning, and DA-W2I and DA-I2W for bi-directional domain adaptation [58]. These modules synergistically enhance generalization on the agricultural target domain, especially under limited label availability, leveraging both image-level and feature-level adaptation strategies.
Similarly, Espejo-García et al. [59] proposed a method that integrates agricultural transfer learning with Generative Adversarial Networks (GANs) to synthesize annotated images, addressing the scarcity of labeled data in agricultural datasets.
Supervised convolutional neural networks (CNNs) have also shown strong performance in image segmentation tasks. In agricultural remote sensing, fully supervised CNNs have been applied to farmland boundary extraction. Lu et al. [60] introduced an attention mechanism to focus on critical spatial features and designed a feature fusion module for enhanced segmentation. However, low-resolution feature maps may cause blurred boundaries due to upsampling. To mitigate this, Li et al. [61] proposed a pixel-wise contextual modeling approach with post-processing to improve boundary clarity.
To further address scale variation issues, Shang et al. [62] presented a multi-scale object extraction framework with attention mechanisms, enabling adaptive segmentation of agricultural objects with diverse sizes. Zhang et al. [63] reformulated farmland extraction as an edge detection problem and developed a High-Resolution Boundary Refinement Network (HBRNet) to improve boundary precision.
Nevertheless, the application of supervised deep DA in real-world agriculture remains constrained by the high cost of acquiring labeled target domain samples. For instance, in the Cityscapes dataset [64], manual annotation of a single natural scene image requires approximately 1.5 hours. This limitation underscores the growing importance of semi-supervised and unsupervised DA paradigms, which are increasingly favored for their ability to leverage unlabeled data and reduce annotation overhead in practical agricultural deployments.
4.2 Semi-supervised Deep Domain Adaptation (DA)
Semi-supervised Deep Domain Adaptation (DA) methods utilize fully labeled samples from the source domain along with a limited amount of labeled or pseudo-labeled data from the target domain. These approaches implement domain transfer through strategies such as self-training, co-training, and hybrid supervision. Their primary advantage lies in striking a balance between annotation cost and model performance, making them particularly valuable in agricultural scenarios characterized by scarce labeled target data.
Self-training and pseudo-label generation are core techniques in semi-supervised DA. Recent innovations have integrated diffusion models into this paradigm, showing promising results in agricultural applications. Ghanbari et al. [65] proposed a framework that combines diffusion models with meta-learning to perform wheat ear segmentation using only a few annotated samples. Specifically, the method requires just three manually labeled wheat images and unlabelled video frames to generate a large pseudo-labeled dataset through a probabilistic diffusion process.
The architecture features a dual-branch encoder-decoder network designed to enable effective feature alignment and semantic consistency between domains (see Figure 9). The model achieved a Dice score of 63.5% on an external test set comprising farmland images from multiple countries, demonstrating strong generalization capabilities. This diffusion-based pseudo-labeling mechanism substantially improves the adaptability of deep DA in data-scarce agricultural environments, offering both scalability and practical relevance.
Figure 9: Schematic of the model architecture in Ghanbari’s method [65].
In this architecture, the encoder learns joint representations for both synthetic and real images, while the mask decoder generates segmentation masks. The image decoder reconstructs real images, encouraging the encoder to adapt to real domain features. Experimental results demonstrate Dice scores of 80.7% on internal test sets and 64.8% on external test sets from 18 plots across five countries. These results highlight the framework’s potential to support generalizable agricultural vision models and facilitate real-world deployment.
By incorporating diffusion models, this approach effectively harnesses large-scale unlabeled datasets and has shown strong performance in image processing tasks [66, 67], underscoring its relevance in precision agriculture.
To address wheat disease detection, Najafian et al. [68] developed the WheatSeedBelt dataset and designed a pipeline combining pretrained models with semi-supervised fine-tuning, achieving high accuracy in Fusarium-damaged kernel (FDK) classification. Subsequently, they proposed a semi-self-supervised wheat head detection method [69], using few annotated images and video sequences to generate pseudo-labeled datasets enhanced by domain adaptation techniques (mAP = 0.827).
Furthermore, they introduced a hybrid approach leveraging synthetic data and multi-domain adaptation for wheat head segmentation [70]. With only two labeled samples, their model achieved a Dice score of 0.89 on internal data and 0.91 after fine-tuning on diverse external data. This greatly enhances model generalization under label-scarce conditions, demonstrating practical viability for real-world deployment.
In vegetable detection, Yang et al. [71] proposed TDA-YOLO, a YOLOv5-based model adapted for tomato detection in densely planted fields under variable lighting. The model employs neural preset color style transfer to generate a pseudo dataset, narrowing the domain gap. Semi-supervised learning and knowledge distillation further enhance target domain adaptability, while the lightweight FasterNet backbone improves inference speed.
Zhu et al. [72] introduced Semi-supervised Center-based Discriminative Adversarial Learning (SCDAL) for domain-adaptive aerial scene classification. Center loss-guided adversarial learning strengthens domain discriminability.
Teng et al. [73] proposed the Classifier-constrained Deep Adversarial Domain Adaptation (CDADA) method, which integrates the Maximum Classifier Discrepancy (MCD) framework with deep convolutional neural networks (DCNNs) to enable semi-supervised cross-domain classification in remote sensing by adversarially aligning feature distributions.
In summary, semi-supervised domain adaptation effectively leverages limited labeled data to guide feature alignment, reduce annotation burdens, and improve model robustness across varied agricultural environments. These approaches provide practical solutions for multimodal data fusion and stage-aware crop monitoring—both essential components of precision agriculture.
4.3 Unsupervised Deep Domain Adaptation (DA)
Unsupervised Deep Domain Adaptation (Unsupervised Deep DA) fully leverages labeled source domain data and unlabeled target domain samples to achieve domain-invariant representation learning through feature space transformation or generative models[74, 75]. These approaches can be categorized into adversarial-based methods, discrepancy-based methods, and self-supervised learning, demonstrating significant value in annotation-scarce scenarios such as agricultural robotics and remote sensing mapping[76].
4.3.1 Adversarial Training
Adversarial learning, particularly via Generative Adversarial Networks (GANs), is widely employed in Unsupervised Domain Adaptation (UDA), primarily for aligning the data distributions of the source and target domains. Two of the most influential adversarial training techniques are the Domain-Adversarial Neural Network (DANN) and the Gradient Reversal Layer (GRL), both of which are conceptually derived from the GAN framework.
GANs, as a class of deep generative models [77], are typically applied in tasks involving image synthesis. Training a GAN requires a dataset representative of the desired target domain, such as CelebA for facial images [78], MNIST for handwritten digits [79], LSUN for bedroom scenes [80], or general object categories like CIFAR-10 [81] and ImageNet [82]. Once trained, the generative model can synthesize images that closely resemble real samples from the training set.
A GAN consists of two neural networks in competition [83]:
  • The generator takes a noise vector as input (typically sampled from a uniform or Gaussian distribution) and attempts to generate synthetic samples that are indistinguishable from real data.
  • The discriminator receives either real samples from the dataset or fake samples from the generator and aims to distinguish between them.
During training, these two networks play a minimax game, in which the generator seeks to fool the discriminator, while the discriminator learns to avoid being deceived. According to Goodfellow et al. [77], this game is defined by the following value function:
$\min_{G}\max_{D}V(D,G)$ $=\mathbb{E}_{x\sim P_{\mathrm{data}}(x)}[\log D(x)]$
$\quad+\mathbb{E}_{z\sim P_{z}(z)}[\log(1-D(G(z)))]$ (6)
Here, $x\sim P_{\text{data}}(x)$ denotes a real sample from the data distribution, $z\sim P_{z}(z)$ is a noise vector sampled from a prior distribution, $D(x;\theta_{d})$ is the discriminator’s probability output for classifying $x$ as real, and $G(z;\theta_{g})$ is the generator’s output given noise $z$.
The objective for the discriminator $D$ is to maximize the probability of correctly classifying real and fake samples, while the generator $G$ aims to minimize the probability that its fake outputs are correctly identified. That is, when $D(G(z))=0$, the discriminator perfectly identifies fake data, and when $D(G(z))=1$, the generator has successfully fooled the discriminator.
Figure 10: Principle diagram of GAN adversarial training. The generator attempts to produce realistic samples, while the discriminator learns to distinguish real from synthetic data.
This adversarial process allows the generator to continuously improve its outputs, ultimately leading to the generation of highly realistic synthetic samples. GANs have proven especially valuable in domain adaptation tasks involving significant visual discrepancies between domains, such as style transfer between different sensors in agricultural remote sensing.
In adversarial training, Generative Adversarial Networks (GANs)—such as CycleGAN—and Discriminative Adversarial Networks (DANN) serve distinct purposes and exhibit different application scenarios. GANs are primarily designed to synthesize data samples, making them especially suitable for tasks such as image style transfer and image generation. In contrast, discriminative adversarial networks focus on aligning feature distributions between domains and are commonly applied in feature representation learning and domain adaptation tasks.
A typical GAN consists of two components: a generator and a discriminator. The generator is trained to produce target domain samples based on the distribution of source domain data, while the discriminator attempts to distinguish real samples from generated ones. In the agricultural domain, GANs have been extensively used for cross-sensor image style transfer. For instance, CycleGAN has been applied to transform the visual style of source domain remote sensing imagery to resemble that of the target sensor, thus mitigating domain shifts [84].
Zhang [85] proposed an unsupervised adversarial domain adaptation method for agricultural land extraction, introducing a multi-scale feature fusion (MSFF) module to adapt across datasets with varying spatial resolutions. This design enhances the model’s robustness in learning domain-invariant features.
Figure 11: Multi-Scale Cross-Domain Adversarial Training and Feature Fusion Framework by Zhang [85].
Ji et al. [86] introduced an end-to-end GAN-based DA framework for land cover classification using multi-source imagery. In this approach, source images are translated into the target style via adversarial learning, and semantic segmentation is performed using a fully convolutional network (FCN). Similarly, Tasar et al. [87] proposed a standard GAN-based DA pipeline for semantic segmentation of very high-resolution (VHR) satellite images and extended it to a multi-source, multi-target, and lifelong learning setting. Wittich et al. [88] designed a Color Map GAN for consistent appearance adaptation in aerial image classification. Liu et al. [89] developed a GAN-based feature extractor for agricultural scene classification that enhances cross-domain classification accuracy by aligning source and target distributions.
While GANs perform exceptionally well in unpaired style transfer tasks, their major drawbacks include susceptibility to mode collapse and high computational cost during training.
On the other hand, Discriminative Adversarial Networks, such as DANN, incorporate a domain classifier that distinguishes between source and target domain features. During training, a gradient reversal layer enables backpropagation to align feature distributions implicitly, thus improving the generalization capacity of the learned model in the target domain. In agriculture, DANN has been widely adopted for domain-invariant feature extraction and alignment.
For example, Yan et al. [90] proposed a Tri-Adversarial Domain Adaptation (TriADA) framework for pixel-level classification in VHR remote sensing images, which incorporates a domain similarity discriminator to learn transferable classifiers. Wang et al. [91] proposed an Adversarial Domain Adaptation (ADDA) approach for cross-device calibration in hyperspectral imaging systems, aligning spectral distributions across devices (e.g., HySpex and Headwall). Their method achieved a 21.5% reduction in RMSE for predicting relative water content (RWC) in maize, highlighting the method’s effectiveness in cross-platform model transfer.
Figure 12: Basic Network Architecture of DANN for Cross-Domain Feature Alignment.
Additional applications of DANN in agricultural image analysis include Bejiga et al. [92], who applied a DANN-based model for large-scale land cover classification using multispectral data. Rahhal et al. [93] designed a multi-source DA strategy using a min-max entropy optimization technique. Elshamli et al. [94] combined denoising autoencoders (DAEs) with DANN to handle DA in multi-temporal and multi-spatial imagery. Mauro et al. [95] integrated self-attention mechanisms into adversarial networks to address domain variance in multi-temporal land cover classification. Mateo et al. [96] proposed a cross-sensor adversarial DA method for cloud detection using Landsat-8 and Proba-V imagery. Ma et al. [97] introduced an Adaptive DANN (ADANN) framework that aligns meteorological features and vegetation indices across U.S. agricultural zones, achieving a 34.1% reduction in yield prediction RMSE.
Discriminative adversarial networks offer the advantage of low annotation dependency while significantly enhancing cross-domain generalization. However, they are sensitive to feature extractor quality and may exhibit convergence instability under large domain gaps.
Adversarial DA strategies, including hybrid GAN-DANN frameworks, are being widely used in remote sensing tasks such as scene classification [84], crop classification [98], road extraction [83], and cloud detection [99]. Nonetheless, the use of fixed distance metrics (e.g., MMD) may not suffice for diverse agricultural imagery datasets.
Recent advances further refined adversarial DA. Hoffman et al. [100] first integrated adversarial training into pixel-level domain adaptation. Tsai et al. [101] proposed AdaptSegNet to reduce domain shifts at the segmentation output level. Vu et al. [102] introduced ADVENT, which applies entropy-based adversarial loss to segmentation outputs. Li et al. [103] proposed a bi-directional learning (BDL) approach to enhance bidirectional feature alignment. Zhang et al. [104] designed DFENet, a domain feature enhancement network that models both channel and spatial dependencies to improve discriminability. Zheng et al. [105] presented TSAN, a two-stage adaptive network for multi-target remote sensing scene classification. Rahhal et al. [93] again applied multi-source adversarial adaptation using entropy-based alignment. Makkar et al. [106] used adversarial learning to extract target-discriminative geospatial features.
Finally, Saito et al. [107] proposed an open-set adversarial DA framework that explicitly separates known and unknown target classes. During training, the feature generator either aligns the sample with known source categories or rejects it as an unknown sample, effectively supporting open-set classification in domain shift scenarios.
4.3.2 Discrepancy-Based Alignment
Discrepancy-based deep domain adaptation (DA) methods aim to reduce the divergence between marginal and/or conditional distributions of source and target domains by incorporating distribution alignment metrics—such as Maximum Mean Discrepancy (MMD)—into deep neural network architectures [108]. These adaptation strategies typically employ additional alignment layers to encourage the learning of domain-invariant, task-relevant representations, as illustrated in Figure 13.
Figure 13: Flowchart of Discrepancy-Based Deep DA Methods.
Long et al. [109] pioneered the Deep Adaptation Network (DAN), which extends convolutional neural networks to domain adaptation by aligning the distributions of learned features across domains. Baktashmotlagh et al. [110] proposed a domain-invariant projection approach using MMD, while Shen et al. [111] introduced a Wasserstein Distance Guided Representation Learning (WDGRL) strategy. Sun et al. [112] extended the CORAL framework by proposing Deep CORAL, which aligns activation correlations in deep layers through nonlinear transformations.
In the context of agriculture, Ferreira [113] proposed an Unsupervised Domain Adaptation framework based on vision Transformers (UDA-SegFormer) to improve crop row and gap detection across farms. By aligning domain-specific features and simplifying pseudo-labels, the model achieved superior generalization performance in challenging environments (e.g., shadows, curved rows), outperforming traditional CNN-based networks such as PSPNet and DeepLabV3+. This work represents the first integration of Transformers with UDA in agricultural settings, enabling label-efficient and robust crop monitoring.
Takahashi et al. [34] enhanced agricultural image recognition by performing category-wise feature space alignment. Building on DAN, Long et al. [114] proposed the Multi-Representation Adaptation Network (MRAN) [115] and the Deep Sub-domain Adaptation Network (DSAN) [116], which leverage sub-domain partitioning and multiple feature representations to address inter-domain variability.
Real-world high-resolution satellite (HSR) images often differ across regions in terms of scale, resolution, and spectral style [117]. To address such heterogeneity, Zhu et al. [118] proposed the Weak Pseudo-supervised Decorrelation Sub-domain Adaptation (WPS-DSA) network for land use classification, which effectively handles spatial and temporal feature distribution shifts.
In the field of cotton boll status recognition, Li [119] introduced NCADA, an unsupervised DA method based on Neighborhood Component Analysis, along with the first benchmark dataset for in-field cotton boll imagery. NCADA outperformed traditional classification methods and achieved competitive accuracy under real-world distribution shifts, paving the way for automated phenotyping in precision agriculture.
Figure 14: Example Images from the In-field Cotton Boll (IFCB) Dataset.
For hyperspectral image (HSI) domain adaptation, Garea et al. [120] introduced TCANet, which incorporates Transfer Component Analysis (TCA) into a deep architecture. Ian et al. [77] developed DDA-Net to enable unsupervised knowledge transfer across HSI datasets. Li et al. [121] proposed a two-stage deep domain adaptation (TDDA) framework: Stage I minimizes MMD-based domain shifts via deep embedding; Stage II employs a spatial-spectral twin network to extract discriminative features. Zhang et al. [122] presented TST-Net, which leverages graph convolutional networks (GCNs) and convolutional layers to encode topological and semantic information for cross-scene classification. Wang et al. [123] introduced a GNN-based DA model that jointly aligns domain and class distributions through CORAL integration. Liang et al. [124] proposed AMF-FSL, an attention-driven few-shot learning framework for HSI classification, which employs target-class alignment, domain attention, and multi-source fusion modules.
In plant pathology, Yan et al. [125] designed a hybrid sub-domain alignment strategy for cross-species plant disease severity recognition. Fuentes et al. [126] developed a domain-adaptive classification model for tomato diseases that accommodates environmental variability. However, these approaches fall short in addressing the complexities of cross-species disease diagnosis. To this end, Wu [127] introduced MSUN, a multi-representation sub-domain alignment network with uncertainty regularization. MSUN effectively handles intra-domain ambiguity, inter-class overlap, and inter-domain heterogeneity, which are critical challenges in plant disease recognition under cross-species settings.
Othman et al. [128] proposed a domain adaptation (DA) network for cross-scene classification, utilizing a two-stage strategy of pretraining and fine-tuning. Their method ensures accurate classification of source domain samples while aligning feature distributions between domains and preserving the geometric structure of the target domain data. Similarly, Lu et al. [129] introduced a Multi-Source Compensation Network (MSCN) designed for cross-scene classification. This architecture incorporates both cross-domain alignment modules and classifier compensation modules, which jointly reduce domain shift and align class semantics across multiple source domains. Zhu et al. [130] developed the Attention-based Multi-Scale Residual Adaptation Network (AMRAN), which combines a residual adaptation module for marginal distribution alignment, an attention module for enhanced feature extraction, and a multi-scale adaptation module that targets conditional distribution alignment. In the SAR image domain, Geng et al. [131] proposed the Deep Joint Distribution Adaptation Network (DJDAN), which simultaneously addresses both marginal and conditional distribution discrepancies for effective transfer learning in SAR image classification tasks.
In recent years, self-supervised learning (SSL) has emerged as a powerful paradigm in the absence of labeled data, demonstrating strong capability in feature representation learning [132]. Zhao et al. [133] proposed a Contrastive Learning and Domain Adaptation Layer (CLA) framework for leaf disease recognition, which innovatively combines self-supervised pretraining with fine-tuned domain adaptation. This two-stage approach is particularly effective in agricultural contexts with large-scale unlabeled image collections.
In the self-supervised pretraining stage, CLA employs contrastive learning to train an encoder on a large pool of unlabeled and distribution-confounded images. The goal is to maximize similarity between augmented views of the same image (positive pairs) while minimizing similarity across different images (negative pairs), akin to the contrastive loss used in SimCLR.
In the second stage, CLA fine-tunes the learned representations with a limited amount of labeled data using a Domain Adaptation Layer (DAL), which jointly optimizes supervised loss and domain alignment loss. The loss function is defined as:
$L_{DAL}=L_{\text{supervised}}+\lambda\cdot L_{\text{MMD}}$ (7)
where $L_{\text{supervised}}$ denotes the conventional classification loss on labeled data, $L_{\text{MMD}}$ represents the domain alignment loss based on Maximum Mean Discrepancy (MMD), and $\lambda$ is a hyperparameter that balances the two objectives.
Experimental results demonstrated that CLA significantly outperformed baseline and comparison methods in both domain alignment and final classification accuracy, achieving a peak accuracy of 90.52%. Further ablation studies and sensitivity analyses were conducted to identify key contributing factors, providing deeper insights into the underlying mechanisms.
The CLA framework provides a compelling example of combining self-supervised contrastive learning with domain adaptation to improve generalization performance in leaf disease recognition under low-label regimes. Its capability to leverage unlabeled data while mitigating domain shifts holds strong potential for broad adoption in agricultural image analysis. By enhancing robustness to domain variance and environmental noise, CLA offers a scalable and label-efficient solution for intelligent plant health monitoring systems.
4.4 Summary of Deep Domain Adaptation (DA) Methods
With the rapid development of deep learning technologies, deep domain adaptation (DA) methods have demonstrated significant advantages in agricultural image analysis. These methods, utilizing deep neural networks (DNNs) and adversarial training, are capable of automatically learning domain-invariant feature representations, effectively addressing the complex problem of domain shift. Typically, deep DA methods are classified into supervised, semi-supervised, and unsupervised categories.
In supervised deep DA methods, source domain classification loss and cross-domain alignment loss are jointly optimized to align the feature distributions between the source and target domains, thereby improving the model’s generalization ability in the target domain. Semi-supervised deep DA methods combine fully labeled data from the source domain with a small amount of labeled data from the target domain, balancing labeling costs with model performance, making them particularly suitable for agricultural scenarios with limited labeled data. Unsupervised deep DA methods rely entirely on labeled data from the source domain and unlabeled data from the target domain, achieving domain adaptation through feature alignment or generative models. These methods have a wide range of applications.
The primary advantage of deep DA methods lies in their ability to handle high-dimensional, complex agricultural data and, through end-to-end learning, automatically extract deep features. This enables them to effectively mitigate domain shifts caused by differences in sensors, environmental changes, and crop growth stages. However, deep DA methods also present certain challenges, including high computational resource demands, long model training times, and poor model interpretability. These issues can be limiting factors in agricultural environments with limited resources or high real-time performance requirements. Nonetheless, with continuous technological advancements, deep DA methods are poised to play an increasingly important role in precision agriculture and smart farming.
5 Experimental Results: Evaluation, Comparison, and Analysis
To effectively demonstrate the superiority of deep domain adaptation (DA) methods in agricultural applications, this section presents experimental results on three core tasks: plant disease detection, crop yield prediction, and agricultural land extraction from remote sensing imagery. Both quantitative metrics and visual results are used to assess model performance.
5.1 Evaluation Metrics
We adopt a diverse set of evaluation metrics to comprehensively assess the performance of classification, regression, and segmentation models in agricultural image analysis. The definitions and corresponding formulas are listed below.
Average Accuracy (AA)
$\text{AA}=\frac{1}{C}\sum_{i=1}^{C}\frac{N_{ii}}{\sum_{j=1}^{C}N_{ij}}$ (8)
Where: $C$ is the number of classes, and $N_{ij}$ denotes the number of samples whose true label is class $i$ and predicted as class $j$. This metric gives equal weight to all classes regardless of class imbalance.
Root Mean Squared Error (RMSE)
$\text{RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}$ (9)
Where: $y_{i}$ is the ground truth value, and $\hat{y}_{i}$ is the predicted value. RMSE expresses prediction error in the same units as the target variable and is widely used in regression tasks.
Coefficient of Determination ($R^{2}$)
$R^{2}=1-\frac{\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}{\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}}$ (10)
Where: $\bar{y}$ denotes the mean of the ground truth values. The $R^{2}$ score measures the proportion of variance in the dependent variable that is predictable from the independent variables.
Intersection over Union (IoU)
$\text{IoU}=\frac{|P\cap G|}{|P\cup G|}$ (11)
Where: $P$ is the predicted region and $G$ is the ground truth region. IoU is a widely used metric for evaluating semantic segmentation performance.
Completeness (COM)
$\text{COM}=\frac{\text{TP}}{\text{TP}+\text{FN}}$ (12)
COM quantifies the proportion of actual agricultural land pixels that were correctly identified.
Correctness (COR)
$\text{COR}=\frac{\text{TP}}{\text{TP}+\text{FP}}$ (13)
COR reflects the proportion of predicted agricultural land pixels that are truly correct.
Segmentation F1 Score
[…]
This F1 score variant harmonizes completeness and correctness to evaluate the quality of segmentation results.
Table 2: Summary of datasets used in the experiments
Dataset Resolution Sensor Image Size Training Set Test Set
DeepGlobe 0.5 m/pixel WorldView-2 2448 $\times$ 2448 30,470 N/A
GID 4 m/pixel GF-2 6800 $\times$ 7200 73,490 734
LoveDA 0.3 m/pixel Spaceborne 1024 $\times$ 1024 15,829 324
PlantVillage 256 $\times$ 256 RGB Fixed size 50,000 5,000
PlantDoc Variable RGB Variable 2,000 500
Plant-Pathology 2048 $\times$ 1365 RGB Fixed size 2,921 730
Corn-Leaf-Diseases 256 $\times$ 256 RGB Fixed size 2,500 500
Tomato-Leaf-Diseases 256 $\times$ 256 RGB Fixed size 4,000 1,000
NASS N/A Tabular data Multi-year regionwise N/A N/A
5.2 Plant Disease Detection Experiments
This section presents the classification, detection, and prediction capabilities of deep domain adaptation (DA) methods in agricultural scenarios. We evaluate plant disease detection, crop yield prediction, and remote sensing-based agricultural land extraction. Since the methods are trained and tested on different datasets, we unify the evaluation by selecting common benchmark datasets. The best-performing results are highlighted in bold to facilitate comparison.
To evaluate domain adaptation for plant disease detection, we conducted four sets of experiments using five benchmark datasets: PlantVillage (PVD) [134], PlantDoc [135], Plant-Pathology [136], Corn-Leaf-Diseases, and Tomato-Leaf-Diseases. Among them, PlantVillage (PVD) was consistently used as the source domain, as it features a controlled acquisition environment with clean backgrounds that facilitates model training.
Figure 15: Domain adaptation setups for four experiments. Left: source domain dataset (PVD); Right: respective target domain datasets.
Table 3 summarizes the results of eight models—Baseline, DAAN, D-CORAL, DAN, DANN, MRAN, DSAN, and MSUN—on the four experiments: C-PD, PVD-PP, PVD-CLD, and PVD-TLD. Compared with the multi-species transfer scenario in C-PD, single-species domain adaptation (e.g., PVD-CLD and PVD-TLD) generally performs better due to clearer intra-class boundaries. MSUN consistently achieves top performance across most tasks, benefiting from its multi-representation modeling and subdomain alignment strategies that address high inter-class similarity and intra-class variability.
Table 3: Accuracy (%) results across four experimental setups: C-PD, PVD-PP, PVD-CLD, and PVD-TLD.
Task Class / Crop Baseline DAAN D-CORAL DAN DANN MRAN DSAN MSUN
C-PD Apple 67.38 68.64 69.06 70.29 69.63 68.98 65.51 71.45
Corn 50.07 48.81 49.11 52.15 48.11 49.87 51.39 56.46
Grape 71.36 63.93 79.79 78.24 83.41 82.38 79.79 85.49
Pepper bell 84.19 87.79 86.75 90.91 90.90 92.46 90.91 93.49
Potato 50.34 63.66 66.06 64.72 62.33 57.82 64.19 67.42
Tomato 30.17 39.82 43.32 46.23 49.51 49.34 45.75 48.67
Average 30.78 33.08 33.37 33.79 39.71 39.84 42.87 56.06
PVD-PP Gray spot 79.54 66.16 77.56 89.02 85.12 81.28 83.96 87.32
Rust 24.10 55.81 33.42 34.99 30.26 52.57 35.42 57.71
Healthy 11.02 36.64 48.05 11.84 28.71 70.87 51.24 73.86
Average 39.27 53.75 52.41 48.89 50.04 67.45 56.41 72.31
PVD-CLD Gray spot 61.49 78.57 79.44 68.29 76.66 79.22 81.71 80.09
Blight 86.75 92.27 91.88 91.50 91.73 92.05 91.58 93.81
Healthy 79.05 84.29 82.64 81.33 82.64 83.61 83.42 84.77
Rust 99.18 100.0 100.0 100.0 100.0 100.0 100.0 100.0
Average 75.96 90.35 89.90 87.89 89.47 91.06 90.69 96.78
PVD-TLD Bacterial spot 5.34 10.23 12.50 12.22 8.24 18.47 10.51 29.27
Healthy 70.61 83.75 85.81 83.41 86.92 75.92 92.03 78.67
Late blight 79.05 3.97 1.64 0.93 3.50 2.11 1.64 3.28
Mold leaf 25.03 33.15 37.43 38.18 38.73 33.15 29.61 32.77
Septoria spot 6.36 5.16 18.73 67.11 4.13 78.64 0.29 80.79
Average 21.46 30.96 32.31 42.80 28.71 44.86 27.57 50.58
In the plant disease transfer detection task, MSUN demonstrates significant advantages across a range of experimental configurations. Whether under multi-plant mixed transfer (C-PD) or single-plant transfer (PVD-PP, PVD-CLD, PVD-TLD), MSUN consistently achieves superior average accuracy compared to mainstream domain adaptation methods.
Specifically, in the C-PD task, MSUN achieves an average accuracy of 56.06%, outperforming the baseline model by more than 25 percentage points. It also surpasses advanced methods such as DSAN and MRAN, illustrating its strong adaptability to large domain shifts and complex inter-class relationships.
In the PVD-PP experiment, MSUN obtains the highest accuracy across all three disease categories, reaching an average accuracy of 72.31%. It significantly outperforms the second-best method MRAN (67.45%) and the traditional DANN (50.04%), reflecting its enhanced transferability in single-plant scenarios.
In the PVD-CLD setup, MSUN achieves an impressive average accuracy of 96.78%, particularly excelling in the “blight” (93.81%) and “healthy” (84.77%) categories. This highlights MSUN’s strong ability to capture discriminative features and reduce inter-class confusion.
The PVD-TLD task, featuring fine-grained categories and highly imbalanced distributions, presents greater challenges. While most methods perform poorly, MSUN still attains an average accuracy of 50.58%. Notably, it achieves 80.79% in the low-sample “Septoria spot” category, demonstrating superior generalization to minority classes and robustness to data imbalance.
Overall, MSUN exhibits consistent and leading performance across various plant disease transfer tasks, validating the effectiveness of its multi-representation learning and subdomain alignment strategy. Compared to traditional domain adaptation approaches, MSUN not only improves classification accuracy but also offers enhanced robustness and generalization in the face of class imbalance, multi-disease complexity, and cross-plant variability.
5.3 Crop Yield Prediction Experiments
For crop yield prediction, we selected the U.S. Corn Belt as the study region, the world’s largest corn-producing area with rich historical yield data. The model training incorporated three data types: remote sensing products, meteorological observations, and crop yield records. All data were sourced from the USDA National Agricultural Statistics Service (NASS, 2020).
The experiment comprised both local (intra-region) and cross-domain (inter-region) prediction tasks. We compared four models: Random Forest (RF), Deep Neural Network (DNN), Domain-Adversarial Neural Network (DANN), and an improved version (ADANN).
Figure 16: Corn planting counties in two ecosystem regions (Eastern Temperate Forest (ETF) and Great Plains (GP)) used in cross-domain prediction experiments.
[…]
Table 4: Comparison of model performance in crop yield prediction across different years and transfer settings.
Year Transfer Setting Metric RF DNN DANN ADANN
[…]
RMSE 1.39 1.18 1.65 1.05
[…]
RMSE 1.13 1.01 1.18 0.92
[…]
RMSE 1.02 1.22 1.21 0.86
[…]
RMSE 1.98 1.82 1.54 1.35
[…]
RMSE 1.34 1.26 1.30 1.24
[…]
RMSE 1.23 0.97 1.17 0.91
[…]
RMSE 1.26 1.36 1.58 1.03
[…]
RMSE 1.78 1.54 1.45 1.23
[…]
RMSE 1.11 1.05 1.00 0.91
[…]
RMSE 0.97 0.95 1.00 0.84
[…]
RMSE 1.01 1.20 1.17 0.97
[…]
RMSE 1.57 1.63 1.14 1.07
[…]
RMSE 1.14 1.15 1.53 1.08
[…]
RMSE 1.09 0.96 1.08 1.05
[…]
RMSE 1.25 1.19 1.11 1.02
[…]
RMSE 2.00 1.80 1.47 1.15
[…]
In addition, inter-annual environmental variability significantly affected model performance. For instance, abnormal wet conditions in parts of the Corn Belt in 2019 weakened the correlation between remote sensing features and actual yields, exacerbating domain discrepancies and degrading traditional model performance. However, ADANN maintained stable and accurate predictions under these extreme conditions, showcasing its strong robustness and adaptability to real-world variations.
Overall, the four-year experimental results indicate that ADANN consistently performs well across both local and cross-domain settings. In particular, its superior accuracy and lower error margins in domain transfer tasks underscore its suitability for cross-region yield modeling. Traditional models, while effective in localized scenarios, exhibit significant performance degradation in cross-domain applications due to their inability to adapt to domain shifts. ADANN, by contrast, emerges as the most effective solution for cross-domain agricultural remote sensing yield prediction, thanks to its stability, generalization ability, and adaptability.
[…]
In summary, ADANN effectively extracts domain -invariant features through adversarial training, mitigating distribution shifts and reducing prediction bias. Its superior performance across four consecutive years in the U.S. Corn Belt validates its practicality and robustness. By maintaining a favorable balance between adversarial domain alignment and prediction error minimization, ADANN delivers stable and accurate cross-domain yield predictions, setting a new benchmark for agricultural remote sensing modeling.
5.4 Remote Sensing-Based Agricultural Land Extraction Experiments
In the comparative experiment for agricultural land extraction from remote sensing images, we selected six representative domain adaptation methods for performance evaluation, including the baseline model (trained solely on the source domain without adaptation), AdaptSegNet [101], ADVENT [102], BDL [103], IntraDA [66], and TransFusion-DualDA [85]. The evaluation was conducted across three benchmark agricultural datasets that vary significantly in spatial resolution and scene complexity:
  • DeepGlobe: High-resolution imagery at 0.5 meters per pixel, containing fine-grained farmland boundary annotations;
  • LoveDA: Ultra-high-resolution (0.3 m/pixel) scenes from urban-rural transition zones with heterogeneous land use;
  • GID: Moderate-resolution (4 m/pixel) dataset covering large regional farmland distributions.
Figure 18: Sample images from three agricultural land datasets. (a) DeepGlobe (0.5 m/pixel); (b) LoveDA (0.3 m/pixel); (c) GID (4 m/pixel).
[…]
Source $\rightarrow$ Target Method IoU COM COR F1
DeepGlobe $\rightarrow$ LoveDA Source-only 36.327 45.895 62.861 49.108
AdaptSegNet 46.921 74.478 55.190 60.404
ADVENT 49.392 74.621 58.085 61.169
BDL 52.234 79.747 59.011 65.334
IntraDA 51.710 82.855 57.341 64.589
TransFusion-DualDA 55.763 81.370 62.549 67.750
GID $\rightarrow$ LoveDA Source-only 36.229 40.139 83.018 46.026
AdaptSegNet 42.931 54.221 71.295 55.454
ADVENT 45.035 60.762 65.021 58.250
BDL 44.592 58.231 67.447 57.732
IntraDA 48.254 60.365 72.898 60.799
TransFusion-DualDA 53.470 92.109 56.891 66.042
DeepGlobe $\rightarrow$ GID Source-only 26.986 39.204 62.884 36.599
AdaptSegNet 43.098 68.614 59.414 56.067
ADVENT 45.995 77.435 57.403 58.940
BDL 46.348 75.789 58.373 59.508
IntraDA 47.631 68.488 60.293 61.465
TransFusion-DualDA 49.553 88.545 54.535 62.439
Table 5: Performance comparison of domain adaptation methods across different dataset transfer settings.
[…]
With ADVENT (d) and IntraDA (e), visual quality improves. ADVENT leverages adversarial learning to mitigate domain gaps, while IntraDA introduces intra-class consistency constraints to improve boundary smoothness. However, both methods still exhibit issues such as boundary blurring and fragmented regions.
TransFusion-DualDA (f) produces the most consistent and precise results, with boundaries closely matching ground truth and accurate region-level predictions. This performance is attributed to its long-range dependency modeling and all-scale feature fusion, which enable simultaneous capture of fine details and global structure. Nonetheless, minor edge artifacts and noise remain, suggesting room for further enhancement in handling high-resolution image heterogeneity.
In conclusion, TransFusion-DualDA demonstrates superior generalization across different spatial resolutions and land types. Its advantages in COM and IoU metrics confirm its robustness in shape-preserving segmentation tasks, making it a powerful tool for unsupervised domain adaptation in remote sensing-based agricultural land monitoring.
5.5 Summary of Experimental Findings
In the plant disease transfer detection task, MSUN demonstrates a leading performance with an average accuracy of 56.06%, surpassing other approaches such as DSAN (42.87%). It particularly excels in the cross-domain classification of corn, grape, sweet pepper, potato, and apple diseases, indicating strong adaptability to complex agricultural scenarios. Although its detection accuracy for tomato diseases is slightly lower (48.67%), models like DANN (49.51%) and MRAN (49.34%) offer potential optimization baselines. Notably, sweet pepper classification yields high accuracy across all methods (above 84%), due to its distinct morphological features.
The superior performance of MSUN stems from its innovative architectural design. Multi-representation modules are employed to extract diverse features, reducing both domain and category discrepancies. Sub-domain adaptation enhances fine-grained semantic alignment, while uncertainty regularization effectively suppresses the noise introduced by pseudo-labels. These integrated mechanisms enable MSUN to achieve optimal results on four public benchmark datasets, positioning it as a preferred solution for multi-crop disease diagnosis with improved generalization and accuracy.
[…]
Model interpretability analysis suggests that ADANN effectively balances domain loss and prediction loss to mitigate the effects of domain shift. Error spatial distribution analysis (Figure 17) highlights that RF and DNN exhibited strong error clustering in regions like Michigan and Ohio. While DANN partially alleviates this through adversarial training, ADANN further optimizes prediction errors in core agricultural zones such as Illinois, underscoring its robustness under dynamic cross-region shifts.
[…]
Overall, the experiments validate the effectiveness of domain-adversarial training and entropy-guided sub-domain partitioning in TransFusion-DualDA, especially under cross-sensor and cross-region conditions. This method significantly reduces the dependency on manual annotations by narrowing the domain gap in complex environments such as urban-rural transitions. Future work will focus on optimizing global feature extraction in high-resolution scenes to further enhance adaptability in dynamic remote sensing applications.
6 Challenges and Future Research Directions
Current research in agricultural intelligence confronts multifaceted challenges, spanning data complexity, methodological adaptability, and technological limitations. While domain adaptation (DA) techniques have shown promise in alleviating cross-domain issues in agricultural image analysis—such as environmental heterogeneity and annotation scarcity—they still face significant bottlenecks in computational efficiency, model interpretability, and generalizability under complex field scenarios. This section highlights key practical pain points and proposes targeted research directions.
6.1 Model Adaptability Barriers in Field Scenarios
The application of deep models in real-world agricultural environments is hindered by the following core issues:
1) Multi-scale spatiotemporal data fusion: There are significant resolution disparities between satellite imagery (10–100 m/pixel), drone imagery (0.1–1 m/pixel), and ground sensor data (centimeter-level). Existing 3D convolutional approaches struggle to achieve effective cross-scale feature alignment.
2) Severe labeled data scarcity: Field datasets often contain fewer than 50 samples per class, especially for rare crop diseases. Traditional transfer learning models exhibit a performance drop exceeding 30% when training sample sizes fall below 100.
3) Limitations in multimodal collaboration: Hyperspectral imaging in field settings suffers from radiometric fluctuations (±30% in light intensity), which induce distributional shifts across time periods. Such variance reduces the effectiveness of direct multimodal integration.
6.2 Key Future Research Directions
To overcome these challenges, four critical research directions are proposed:
(1) Lightweight Spatiotemporal Modeling Framework: Develop ultra-compact network architectures with fewer than 1M parameters, incorporating physical priors (e.g., crop growth models) to guide feature extraction. Emphasize spatiotemporal decoupling to enable cross-scale fusion between drone (5 cm) and satellite (10 m) imagery. Aim to reduce computational energy consumption to 20% of current 3D CNN baselines.
(2) Weakly Supervised Transfer Learning Paradigm: Construct pseudo-labeling mechanisms leveraging temporal continuity in crop growth stages to generate self-supervised learning signals. Implement curriculum learning strategies informed by phenological patterns to maintain over 85% classification accuracy with under 100 labeled samples.
(3) Robust Multimodal Fusion System: Design spectral normalization-based feature extraction to suppress illumination-related noise. Develop physically informed adversarial training to ensure cross-modality consistency under ±40% lighting changes. Address cross-view spatial alignment between aerial and ground sensors by establishing sub-meter correspondence standards.
(4) Privacy-Preserving Collaborative Learning: Enhance communication efficiency in federated learning frameworks to support daily updates within 10 Mbps bandwidth. Introduce differential privacy strategies compliant with GDPR, maintaining performance degradation below 5% with a privacy budget of $\epsilon=2$.
To support real-world deployment, a standardized open benchmark is recommended, including over 10 typical crops across three or more climate zones. The benchmark should incorporate field-specific noise factors (e.g., device error ±5%, occlusion rate 20%) to ensure alignment between algorithmic development and application robustness.
7 Conclusion
Adaptive modeling in agriculture faces persistent challenges, particularly in handling the complexity of spatiotemporal dynamics, the scarcity of annotated samples, and the heterogeneity of multimodal data sources. As agricultural environments grow more complex, the effectiveness of existing domain adaptation techniques in field scenarios—especially those involving cross-seasonal, cross-regional, and cross-modal variability—remains limited.
Future research should focus on building task-specific high-dimensional adaptive models to enhance performance in diverse agricultural applications, such as crop health monitoring, yield forecasting, and pest recognition. Unsupervised domain adaptation is especially valuable under annotation-scarce conditions, while multimodal fusion —including image, weather, and soil data—holds potential for further improving model generalization. Moreover, source-free domain adaptation represents a critical frontier where access to source data is restricted due to privacy or operational constraints.
The field-deployed robustness of these models hinges on their ability to operate under unpredictable environmental conditions, diverse crop types, and varying sensor platforms. Addressing these technical and operational barriers is essential for advancing precision agriculture and intelligent farming systems. Improved cross-domain adaptability and model resilience will ultimately support more accurate agricultural decision-making, yield estimation, and disease surveillance, contributing to the transformation of modern agricultural practices.
Acknowledgment
This project is jointly supported by the High-Quality Development Special Project of the Ministry of Industry and Information Technology (TC240A9ED-56), the Academician Workstation Program of Yunnan Province (202405AF140013), the Shanghai Agricultural Technology Innovation Project (2024-02-08-00-12-F00032), the Science and Technology Program of Qinghai Province (2024-NK-141S-4), and the Shanghai Municipal Commission of Science and Technology (Grant No. 24TS1415900).